# Particles before symmetry

### 05 Mar 2026 v5

@@ 3.2 Dynamical Mass Generation for the Higgs in the Linearised Theory
We assume $E$ admits at least one unit section and pick one, $e_{0}(x)$. Again, two different choices $e_{0},e^{\prime}_{0}$ will be isomorphic: they are structurally indistinguishable within (their copies of) $E$, since they have the same (unit) norm (and there is no more background structure on $E$). Thus, in the usual terminology, any two such choices can be mapped to each other by an element of $\mathsf{Aut}(E)$. So, in order to parametrise the space of non-isomorphic unit sections, we can fix the same $e_{0}$ across configurations. In the jargon of \@BBOPcitep\@BAP\@BBN(Gomes, 2025a; Kabel et al., 2025; Gomes and Butterfield, 2023)\@BBCP, we take $e_{0}$ to define a ‘representational scheme’ across physical possibilities, so that each configuration $\varphi\in\Gamma(E)$ is parallel to the same $e_{0}$. This is this demonstration’s analogue of choosing ‘unitary gauge’, and it is the only place where symmetry could be mentioned. [^16] Indeed, representational schemes can be compared to gauge-fixings (cf. \@BBOPcite\@BAP\@BBN(Gomes, 2025a, Sec. 3.3)\@BBCP), as follows. Consider $\Gamma(E)$, and its sector $\Gamma_{0}(E)$. Let $\varphi,\varphi^{\prime}\in\Gamma_{0}(E)$. The group $\mathsf{Aut}(E)$ acts transitively on the unit normal sections: it can take any internal direction into any other. Therefore, we could, by a suitable gauge transformation on $\varphi$, make it collinear with $\varphi^{\prime}{}$. Once they are collinear, it is a trivial matter to separate out the part that has a given norm, be it $e_{0}$, $v^{\prime}e_{0}$, or $ve_{0}$ (see below, for the definition of $v$).
@@
where here $V:\Gamma(E)\rightarrow\mathbb{R}$ is assumed to be homogeneous on $M$ and depend only on $\|\varphi\|$, and so $w_{o}(x)=v\,e_{0}(x)\in\Gamma_{0}(E)$ with $\|w_{o}(x)\|=v\neq 0,\,\forall x\in M$. So any potential whose minimum is non-zero suffices for identifying $w_{o}(x)=v\,e_{0}(x)$ for some $v>0.$ The familiar ’Mexi  [^18] Usually, one takes the Higgs vacuum vector $w_{o}$ to break the symmetry, and then one expands the Higgs fields around this vacuum. But then one must find a way to get rid of the ‘Goldstone modes’: directions of the Higgs field that are parallel to the gauge orbit of the vacuum. This is where one invokes unitary gauge. But this is awkward: angles are preserved by gauge transformations, so how could a gauge transformation act to eliminate modes that are defined by their inner product with respect to the Higgs vacuum, which is assumed to be a physical object?
@@
The mass terms will be obtained from (3.16) by diagonalising $\mathsf{Hess}(V)_{w_{o}}$. Thus, as in the standard account, the mass terms for different components of the Higgs along the different directions within $T_{w_{0}}\Gamma_{0}(E)$ are given by the eigenvalues of the Hessian with those directions as eigenvectors. But, per (3.10), in the current case we are considering only $\delta\varphi=H(x)e_{0}\in T_{w_{o}}\Gamma_{0}(E)$: only $e_{0}$ can have non-vanishing eigenvectors (the eigenvalues of any completion of $e_{0}$ into an orthogonal basis for $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 14) would vanish for $e_{I},I\neq 0$). The value of the surviving eigenvalue along $e_{0}$ will depend on the form of $V(\|\varphi\|)$, which is assumed to be translationally invariant (or homogeneous on $M$). Here we will just call this eigenvalue $m_{H}^{2}$. [^19] In the standard case where $V=-\mu\|\varphi\|^{2}+\lambda\|\varphi\|^{4}$, using (3.9) and (3.10) we can replace $\|\varphi\|^{2}\rightarrow(H+v)^{2}$, and $\frac{\delta}{\delta\varphi}\rightarrow\frac{\partial}{\partial H}$. So $\mathsf{Hess}(V)\rightarrow\frac{\partial^{2}V}{\partial H^{2}}$. Then an easy computation reveals: $\frac{\partial V}{\partial H}=-2\mu(H+v)+4\lambda(H+v)^{3}$ (3.17) and $\frac{\partial^{2}V}{\partial H^{2}}=-2\mu+12\lambda(H+v)^{2},$ (3.18) Setting $\frac{\partial V}{\partial H}=0$ and $H=0$ in (3.17) we obtain $v=\sqrt{\frac{\mu}{2\lambda}}$. Setting $v=\sqrt{\frac{\mu}{2\lambda}}$ and $H=0$ in (3.18) we obtain $\frac{\partial^{2}V}{\partial H^{2}}(w_{o})=4\mu=m_{H}^{2}.$
@@
This concludes the geometric gloss on the Higgs mechanism. No explicit mention of stabilisers, decomposition of gauge orbits, Killing forms on the Lie algebra, Goldstone’s theorem, etc was required—devices indispensable in the standard formulation (see \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8.1)\@BBCP; \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2.2)\@BBCP; \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP for comparison). [^20] For instance, the disappearance of Goldstone modes through a choice of gauge is replaced here by having all $\varphi\in\Gamma_{0}(E)$ parallel to $e_{0}$ and the orthogonality relation (3.4). This is just what you would expect from the geometry alone.
@@ 4.2 The geometry-first presentation
whereas vector bosons are replaced by the corresponding affine covariant derivatives, e.g. $\nabla^{1},\nabla^{2},\nabla^{3}$ (see \@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025b)\@BBCP for more details).
In this formulation, weak isospin $T_{3}$—defined only relative to a chosen basis of the Lie algebra—has no independent geometrical meaning (see Section B.2 and footnote 40). Left-handed fermions are best understood as components of vector fields $\mathbf{Q}_{L}$ and $\mathbf{L}_{L}$. The distinction between electron and electron-neutrino, or between up- and down-left quarks, appears only through their couplings to the Higgs. The Higgs field $\boldsymbol{\varphi}$ provides a frame in $\mathbb{C}^{2}$ that gives $T_{3}$ physical significance. The charges in Table 1 presuppose the frame $\varphi=(0,\varphi^{0})^{T}$; only in that frame do left-handed up-quark components appear as $(u^{I}_{L},0)^{T}$. [^21] This is why Table 1 can mislead: if both Higgs components are retained, the ‘up’ and ‘down’ labels have no physical meaning yet.
@@ 5.1 Slack between symmetry and geometry: examples and lessons
From the PFB-POV this is perfectly admissible: one may posit a group that preserves the relevant geometric structures without exhausting them. Still, there is a geometric interpretation: the action rotates the complex volume form of $\mathbb{C}^{3}$. [^30] That is, it rotates $\Lambda^{3}\mathbb{C}^{3}$ (equivalently, the determinant line). Fixing a unit complex volume form reduces $U(3)$ to $SU(3)$; the residual $U(1)$ is the phase on the determinant line. See footnote 25. Thus in this case we have
@@
so both conditions in (5.2) fail. Intrinsically, $G$ may be either larger or smaller than $\mathsf{Aut}(V)$, and the group can neither be recovered from, nor recover, the geometry of the fibre.
@@
  • The adjoint bundle is $\mathsf{End}_{0}(E):=\{A\in\mathsf{End}(E)\mid\mathsf{Tr}(A)=0,\,A^{\dagger}=-A\}$, the bundle of traceless anti-Hermitian endomorphisms. Its sections are the “gluon fields” in the sense that $\nabla={\mathrm{d}}+\omega$ for $\omega\in\Gamma(T^{*}M\otimes\mathsf{End}_{0}(E))$.
@@
    $G=SU(n)\times SU(m)$ with $\rho$ the projection onto the first factor: the second factor acts trivially on $E$, and cannot be recovered from $\mathsf{Aut}(E_{x})$—slack in the sense of Example 2.
@@
The VB-POV rules out all three by fiat: it insists that the gauge group be the automorphism group of the fundamental bundle. But each of these is a perfectly consistent principal-bundle formulation, and each induces a well-defined covariant derivative on the associated bundle. Yet in each case the group $G$ floats free of the geometry of the fibre: it is either too large, too small, or simply different.
@@
In the PFB-POV, this is perfectly consistent. The connection $\varpi$ carries physical observables independently of matter. The Yang–Mills equations are valued in $\mathfrak{su}(n)$, not $\mathfrak{u}(n)$; the gauge sector has different degrees of freedom. In vacuum Yang–Mills theory, the connection $A_{\mu}$ is a physical field in its own right, with dynamics governed by $G$. If $G=U(n)$ rather than $SU(n)$, the vacuum theory has an extra $U(1)$ factor. These are different theories, with different physical content. Moreover, any Lie-algebra directions in $\ker\rho_{*}$ remain unsourced by matter: the matter current $J(A,\Phi)$ takes values in $\mathrm{im}(\rho_{*})$, so the slack directions satisfy vacuum Yang–Mills equations regardless of the matter configuration (cf. Equation (A.11) in Appendix A).
@@
But that is not our world—at least, not our Standard Model; we do not observe such a mismatch. To be clear: this is not a complaint that practitioners of the PFB-POV secretly believe the principal connection $\varpi$ is a physical entity—though most attribute physical significance to its spacetime representatives $A_{\mu}$, or to a section of the affine bundle of connections \@BBOPcitep\@BAP\@BBN(Jacobs, 2023; Kobayaschi, 1957)\@BBCP; even in vacuum Yang–Mills theory where no matter fields are present. The complaint is that the PFB-POV’s representational resources outrun its explanatory needs. This freedom could in principle be observable—as the slack world illustrates—but in the theories physicists actually take seriously, it is not. I return to this point in Section 5.4.
@@ 5.3 Recovery of the gauge group from matter bundles
Section 5.1 showed that the gauge group can float free of the geometry of any single matter bundle. But the Standard Model has many particle species, each in its own associated bundle. One might hope that, taken together, they pin down $G$: that the slack disappears when the full matter content is taken into account. If so, positing $G$ independently would be a convenience, not a substantive addition. But the hope fails. To recover the product group from the totality of matter bundles, one must also specify which (factors of which) bundles are fundamental—precisely what the VB-POV supplies at the outset.
In more detail: in many gauge theories—and routinely in the Standard Model—the fibres of associated bundles take the form of direct sums or tensor products, such as
$V=V_{1}\otimes V_{2},$ (5.10)
with different factors of the gauge group acting on different components, and sometimes trivially on them. (This is what happens, for instance, when, in the Standard Model, a particle’s representation labels include $\mathbf{1}$ in either of the first two entries or 0 in the last, corresponding to hypercharge). As a result, kernels appear factorwise, and for any given multiplet we have both conditions of (5.2) failing. The Standard Model abounds with such cases: for the right-handed electron bundle, $SU(3)$ and $SU(2)$ act trivially, so no single associated bundle for $e_{R}$ can recover the full $SU(3)\times SU(2)\times U(1)$; similar things can be said about right-handed quarks, etc. In short, the representation spaces defined by the particles’ quantum numbers admit group actions in which entire factors of $G$ play no role—precisely the kind of degeneration illustrated in Eq. (5.5).
From the PFB-POV this is no defect: the standard strategy assigns each particle type to a section of some associated bundle, without attempting to reconstruct the full gauge group from the geometry of any single fibre. This slack poses no internal difficulty, for the two ingredients—the symmetry and the vector bundle structure—are jointly posited and mutually consistent. [^32] If you think that is not how we actually use the principal-associated bundle formalism, I agree; see Section 5.4 below. But if one tries to recover the group from the automorphisms of those associated bundles, the recovery is not merely incomplete; in general, it is inconsistent.
To see this, suppose we are given a collection of associated vector bundles and then want to recover the structure group from the automorphism groups of the associated bundles. Given the collection
$(P,M,G,\{\rho_{i}\}_{i},\{V_{i}\}_{i}),\quad i=1,2$ (5.11)
we can try to reconstruct $(P,G,M)$ from each $L_{\text{\tiny{adm}}_{i}}(E_{i})$ (the bundle of admissible frames for $E_{i}$, see Section 2.1). The problem is that we would in general recover different groups for different $i$: constructing the bundle of admissible frames forces a subgroup $G^{\prime}\subset G\simeq GL(V)$ to act trivially on some subspaces of $E_{i}$, so the resulting principal bundle reflects only the subspaces where $G^{\prime}$ acts non-trivially—and these differ across the associated bundles.
For concreteness, take $V\simeq{\mathbb{C}}^{3}\otimes{\mathbb{C}}^{2}$, $G=SU(3)\times SU(2)$ and consider two representations: $\rho_{1}=\mathbf{3}\otimes\mathbf{1}$ (a colour triplet, singlet under $SU(2)$), and $\rho_{2}=\mathbf{1}\otimes\mathbf{2}$ (a weak doublet, singlet under $SU(3)$). Thus in this case we would recover $G^{\prime}_{1}=SU(3)$ from one bundle, and $G^{\prime}_{2}=SU(2)$ from the other. What we would like, of course, is $SU(3)\times SU(2)$; but the product structure is nowhere to be found at the level of any single associated bundle.
One might hope to do better by considering the entire matter content at once. Suppose we form the direct sum of all the associated bundles,
$E_{\text{tot}}=E_{1}\oplus E_{2}\oplus\cdots\oplus E_{k},$ (5.12)
and try to recover $G$ from the structure of $E_{\text{tot}}$. There are two natural strategies; both fail.
The first strategy examines the full automorphism group $\mathsf{Aut}(E_{\text{tot}})$. This is too large: it includes not only block-diagonal transformations that act independently on each summand, but also off-diagonal maps that mix summands—transformations with no counterpart in gauge theory.
The second strategy restricts to block-diagonal automorphisms: the subgroup of $\mathsf{Aut}(E_{\text{tot}})$ that preserves each summand $E_{i}$ separately. Call this $\prod_{i}\mathsf{Aut}(E_{i})$. Assuming $G$ acts faithfully on the total matter content (as it does in the Standard Model up to the $\mathbb{Z}_{6}$ kernel), $G$ embeds injectively into $\prod_{i}\mathsf{Aut}(E_{i})$, so no factor of $G$ is invisible. But knowing that $G$ lives inside this product does not tell you which subgroup it is. And crucially, the product decomposition that the block-diagonal strategy hands you is indexed by particle species labeled by ii—one $i$ per associated bundle—whereas the gauge group is a product indexed by interactions: $SU(3)$ for colour, $SU(2)$ for weak isospin, $U(1)$ for hypercharge. These are entirely different decompositions. The species-indexed product $\prod_{i}\mathsf{Aut}(E_{i})$ contains the interaction-indexed product $G=G_{1}\times G_{2}\times G_{3}$ as a subgroup, but does not single it out.
To illustrate: a left-handed quark lives in $E_{Q}=\mathbb{C}^{3}\otimes\mathbb{C}^{2}\otimes\mathbb{C}$, with $\mathsf{Aut}(E_{Q})\simeq U(6)$ (preserving the Hermitian structure on the six-dimensional total fibre). The gauge group $SU(3)\times SU(2)\times U(1)$ embeds in $U(6)$, but so do many other groups. Nothing intrinsic to $E_{Q}$ singles out the tensor-product decomposition $\mathbb{C}^{3}\otimes\mathbb{C}^{2}\otimes\mathbb{C}$ from which the product structure of $G$ derives. That decomposition is additional data—it is the specification of fundamental bundles.
Claim. Given a finite collection of Hermitian vector bundles $\{(E_{i},\langle\cdot,\cdot\rangle_{i})\}_{i=1}^{k}$, or even the enriched collection $\{(E_{i},\langle\cdot,\cdot\rangle_{i},\nabla^{i})\}_{i=1}^{k}$ of bundles equipped with compatible covariant derivatives, the product decomposition of the gauge group $G=\prod_{\alpha}G_{\alpha}$ into independent interaction sectors is not uniquely determined—even when $G$ acts faithfully on the total matter content.

### 23 Mar 2026 v6

@@ 3.2 Dynamical Mass Generation for the Higgs in the Linearised Theory
We assume $E$ admits at least one unit section and pick one, $e_{0}(x)$. Again, two different choices $e_{0},e^{\prime}_{0}$ will be isomorphic: they are structurally indistinguishable within (their copies of) $E$, since they have the same (unit) norm (and there is no more background structure on $E$). Thus, in the usual terminology, any two such choices can be mapped to each other by an element of $\mathsf{Aut}(E)$. So, in order to parametrise the space of non-isomorphic unit sections, we can fix the same $e_{0}$ across configurations. In the jargon of \@BBOPcitep\@BAP\@BBN(Gomes, 2025b; Kabel et al., 2025; Gomes & Butterfield, 2023)\@BBCP, we take $e_{0}$ to define a ‘representational scheme’ across physical possibilities, so that each configuration $\varphi\in\Gamma(E)$ is parallel to the same $e_{0}$. This is this demonstration’s analogue of choosing ‘unitary gauge’, and it is the only place where symmetry could be mentioned. [^13] Indeed, representational schemes can be compared to gauge-fixings (cf. \@BBOPcite\@BAP\@BBN(Gomes, 2025b, Sec. 3.3)\@BBCP), as follows. Consider $\Gamma(E)$, and its sector $\Gamma_{0}(E)$. Let $\varphi,\varphi^{\prime}\in\Gamma_{0}(E)$. The group $\mathsf{Aut}(E)$ acts transitively on the unit normal sections: it can take any internal direction into any other. Therefore, we could, by a suitable gauge transformation on $\varphi$, make it collinear with $\varphi^{\prime}{}$. Once they are collinear, it is a trivial matter to separate out the part that has a given norm, be it $e_{0}$, $v^{\prime}e_{0}$, or $ve_{0}$ (see below, for the definition of $v$).
@@
where here $V:\Gamma(E)\rightarrow\mathbb{R}$ is assumed to be homogeneous on $M$ and depend only on $\|\varphi\|$, and so $w_{o}(x)=v\,e_{0}(x)\in\Gamma_{0}(E)$ with $\|w_{o}(x)\|=v\neq 0,\,\forall x\in M$. So any potential whose minimum is non-zero suffices for identifying $w_{o}(x)=v\,e_{0}(x)$ for some $v>0.$ The familiar ‘Mexican hat’ potential $V=-\mu\|\varphi\|^{2}+\lambda\|\varphi\|^{4}$ is one such choice. [^15] Usually, one takes the Higgs vacuum vector $w_{o}$ to break the symmetry, and then one expands the Higgs fields around this vacuum. But then one must find a way to get rid of the ‘Goldstone modes’: directions of the Higgs field that are parallel to the gauge orbit of the vacuum. This is where one invokes unitary gauge. But this is awkward: angles are preserved by gauge transformations, so how could a gauge transformation act to eliminate modes that are defined by their inner product with respect to the Higgs vacuum, which is assumed to be a physical object?
@@
The mass terms will be obtained from (3.16) by diagonalising $\mathsf{Hess}(V)_{w_{o}}$. Thus, as in the standard account, the mass terms for different components of the Higgs along the different directions within $T_{w_{0}}\Gamma_{0}(E)$ are given by the eigenvalues of the Hessian with those directions as eigenvectors. But, per (3.10), in the current case we are considering only $\delta\varphi=H(x)e_{0}\in T_{w_{o}}\Gamma_{0}(E)$: only $e_{0}$ can have non-vanishing eigenvectors (the eigenvalues of any completion of $e_{0}$ into an orthogonal basis for $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 11) would vanish for $e_{I},I\neq 0$). The value of the surviving eigenvalue along $e_{0}$ will depend on the form of $V(\|\varphi\|)$, which is assumed to be translationally invariant (or homogeneous on $M$). Here we will just call this eigenvalue $m_{H}^{2}$. [^16] In the standard case where $V=-\mu\|\varphi\|^{2}+\lambda\|\varphi\|^{4}$, using (3.9) and (3.10) we can replace $\|\varphi\|^{2}\rightarrow(H+v)^{2}$, and $\frac{\delta}{\delta\varphi}\rightarrow\frac{\partial}{\partial H}$. So $\mathsf{Hess}(V)\rightarrow\frac{\partial^{2}V}{\partial H^{2}}$. Then an easy computation reveals: $\frac{\partial V}{\partial H}=-2\mu(H+v)+4\lambda(H+v)^{3}$ (3.17) and $\frac{\partial^{2}V}{\partial H^{2}}=-2\mu+12\lambda(H+v)^{2},$ (3.18) Setting $\frac{\partial V}{\partial H}=0$ and $H=0$ in (3.17) we obtain $v=\sqrt{\frac{\mu}{2\lambda}}$. Setting $v=\sqrt{\frac{\mu}{2\lambda}}$ and $H=0$ in (3.18) we obtain $\frac{\partial^{2}V}{\partial H^{2}}(w_{o})=4\mu=m_{H}^{2}.$
@@
This concludes the geometric account of the Higgs mechanism. No explicit mention of stabilisers, decomposition of gauge orbits, Killing forms on the Lie algebra, Goldstone’s theorem, etc was required—devices indispensable in the standard formulation (see \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8.1)\@BBCP; \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2.2)\@BBCP; \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP for comparison). [^17] For instance, the disappearance of Goldstone modes through a choice of gauge is replaced here by having all $\varphi\in\Gamma_{0}(E)$ parallel to $e_{0}$ and the orthogonality relation (3.4). This is just what you would expect from the geometry alone.
@@ 4.2 The geometry-first presentation
whereas vector bosons are replaced by the corresponding affine covariant derivatives, e.g. $\nabla^{1},\nabla^{2},\nabla^{3}$ (see \@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025a)\@BBCP for more details).
In this formulation, weak isospin $T_{3}$—defined only relative to a chosen basis of the Lie algebra—has no independent geometrical meaning (see Section B.2 and footnote 38). Left-handed fermions are best understood as components of vector fields $\mathbf{Q}_{L}$ and $\mathbf{L}_{L}$. The distinction between electron and electron-neutrino, or between up- and down-left quarks, appears only through their couplings to the Higgs. The Higgs field $\boldsymbol{\varphi}$ provides a frame in $\mathbb{C}^{2}$ that gives $T_{3}$ physical significance. The charges in Table 1 presuppose the frame $\varphi=(0,\varphi^{0})^{T}$; only in that frame do left-handed up-quark components appear as $(u^{I}_{L},0)^{T}$. [^18] This is why Table 1 can mislead: if both Higgs components are retained, the ‘up’ and ‘down’ labels have no physical meaning yet.
@@ 5.1 Slack between symmetry and geometry: examples and lessons
From the PFB-POV this is perfectly admissible: one may posit a group that preserves the relevant geometric structures without exhausting them. Still, there is a geometric interpretation: the action rotates the complex volume form of $\mathbb{C}^{3}$. [^27] That is, it rotates $\Lambda^{3}\mathbb{C}^{3}$ (equivalently, the determinant line). Fixing a unit complex volume form reduces $U(3)$ to $SU(3)$; the residual $U(1)$ is the phase on the determinant line. See footnote 22. Thus in this case we have
@@
so both conditions in (5.2) fail with respect to the single bundle $V$. Note, however, that this form of slack is not an obstacle to VB-presentability: the VB-POV can accommodate the trivially-acting factor $SU(n)$ by introducing a separate fundamental bundle whose automorphism group is $SU(n)$, with no matter field living there. The corresponding gauge sector would satisfy vacuum Yang–Mills equations (cf. Equation (A.14) in Appendix A). What this example shows is that a single matter bundle cannot recover the full gauge group—a point that will matter in Section 5.3.
@@
  • The adjoint bundle is $\mathsf{End}_{0}(E):=\{A\in\mathsf{End}(E)\mid\mathsf{Tr}(A)=0,\,A^{\dagger}=-A\}$, the bundle of traceless anti-Hermitian endomorphisms. Sections of $T^{*}M\otimes\mathsf{End}_{0}(E)$ parametrise the affine space of compatible covariant derivatives: $\nabla=\nabla_{0}+\omega$ for any choice of reference derivative $\nabla_{0}$. The fundamental dynamical object is $\nabla$ itself; $\omega$ is not an independent field but a difference between two derivatives, and appears only upon choosing $\nabla_{0}$. What are called “gluon fields” in the standard formalism correspond, in the VB-POV, to the covariant derivative $\nabla$ on $E$.
@@
    $G=SU(n)\times SU(m)$ with $\rho$ the projection onto the first factor: the second factor acts trivially on $E$, and cannot be recovered from $\mathsf{Aut}(E_{x})$—slack in the sense of Example 2. (As noted there, the VB-POV can absorb this slack by introducing a separate fundamental bundle for $SU(m)$; what is at issue here is that the PFB-POV tolerates the mismatch at the level of any single associated bundle.)
@@
The VB-POV does not tolerate any of these mismatches at the level of a single associated bundle: it insists that each factor of the gauge group be the automorphism group of some fundamental bundle. But it handles them differently. Case (2)—a trivially-acting factor—is absorbable: one introduces a separate fundamental bundle for the idle factor, running vacuum Yang–Mills. Cases (1) and (3), by contrast, involve a mismatch between $G$ and $\mathsf{Aut}(V)$ for the bundle on which $G$ acts nontrivially, and these the VB-POV genuinely excludes.
@@
In the PFB-POV, this is perfectly consistent. The connection $\varpi$ carries physical observables independently of matter. The Yang–Mills equations are valued in $\mathfrak{su}(n)$, not $\mathfrak{u}(n)$; the gauge sector has different degrees of freedom. In vacuum Yang–Mills theory, the connection $A_{\mu}$ is a physical field in its own right, with dynamics governed by $G$. If $G=U(n)$ rather than $SU(n)$, the vacuum theory has an extra $U(1)$ factor. These are different theories, with different physical content. Moreover, any Lie algebra directions in $\ker\rho_{*}$ remain unsourced by matter: the matter current $J(A,\Phi)$ takes values in $\mathrm{im}(\rho_{*})$, so the slack directions satisfy vacuum Yang–Mills equations regardless of the matter configuration (cf. Equation (A.11) in Appendix A).
@@
A trivially-acting factor $H\subset G$ can, however, be absorbed into the VB-POV: one introduces a separate fundamental bundle whose automorphism group is $H$, with no matter field living in it, and the corresponding gauge sector satisfies vacuum Yang–Mills equations. So continuous slack of the kind exhibited in Example 2 is not an obstacle to VB-presentability—a fundamental bundle need not contribute to any matter bundle \@BBOPcitep\@BAP\@BBN(Gomes, 2026)\@BBCP.
The genuine non-equivalence lies elsewhere: in the kernel-entanglement obstruction, in the non-recovery of tensorial genealogy, in the strict inclusion of theory-spaces, and in a difference of symmetry content (outer automorphisms of the gauge group are legitimate PFB-POV symmetries but have no VB-POV counterpart). These obstructions are analysed in \@BBOPcitet\@BAP\@BBNGomes (2026)\@BBCP. The slack world illustrates how the PFB-POV’s representational resources can outrun its explanatory needs; the VB-POV’s response is to structure the slack rather than forbid it. I return to this in Section 5.4.
@@ 5.3 Recovery of the gauge group from matter bundles
Section 5.1 showed that the gauge group can float free of the geometry of any single matter bundle. But the Standard Model has many particle species, each in its own associated bundle. One might hope that, taken together, they pin down $G$: that the slack disappears when the full matter content is taken into account. If so, positing $G$ independently would be a convenience, not a substantive addition. But the hope fails. To recover the product group from the totality of matter bundles, one must also specify which (factors of which) bundles are fundamental—precisely what the VB-POV supplies at the outset. (The arguments of this subsection are developed more formally in \@BBOPcitet\@BAP\@BBNGomes (2026)\@BBCP, where the failure of recovery is given a categorical formulation.)
The difficulty is threefold. First, no single matter bundle sees the full group: for the right-handed electron, $SU(3)$ and $SU(2)$ both act trivially, so the automorphism group of its fibre recovers only $U(1)$. Second, even taking the entire matter content at once does not help. The full automorphism group $\mathsf{Aut}(E_{\text{tot}})$ of the direct sum of all matter bundles is too large (it includes species-mixing maps with no gauge-theoretic counterpart), and restricting to block-diagonal automorphisms yields a product indexed by particle species rather than by interactions—an entirely different decomposition from $G=G_{1}\times G_{2}\times G_{3}$. A left-handed quark, for instance, lives in $E_{Q}\simeq\mathbb{C}^{3}\otimes\mathbb{C}^{2}\otimes\mathbb{C}$, whose Hermitian automorphism group is $U(6)$; nothing intrinsic to this six-dimensional fibre singles out the tensor-product decomposition from which the product structure of $G$ derives. [^29] If you think that is not how we actually use the principal-associated bundle formalism, I agree; see Section 5.4 below. Third, requiring the gauge group to act faithfully on all matter does not resolve the problem—it can make things worse. In the Standard Model, a $\mathbb{Z}_{6}\subset SU(3)\times SU(2)\times U(1)$ acts trivially on every multiplet, and quotienting by it yields a group $G’=(SU(3)\times SU(2)\times U(1))/\mathbb{Z}_{6}$ that is not a product of classical groups at all. This is the kernel-entanglement obstruction: the $\mathbb{Z}_{6}$ sits diagonally across all three factors, so the quotient cannot be expressed as $\prod_{a}G’_{a}$, and no geometry-first starting point can produce it. (These arguments are developed formally in \@BBOPcitet\@BAP\@BBNGomes (2026)\@BBCP, where the failure of recovery is given a categorical formulation: the natural functor from geometry-first data to principal bundles with matter is faithful but neither essentially surjective nor full.)